OpenStudy (anonymous):

(2x3+13x2+16x+5)÷(x+5)

OpenStudy (timaashorty):

\[\frac{ 2x3+13x2+16x+5 }{x+5 } = \] \[(2x+1) (x+1) =\] \[x(2x+3)+1=\] \[2x ^{2} +3x+1\]

OpenStudy (timaashorty):

& welcome to openstudy (:

OpenStudy (anonymous):

THANK YOU SO MUCH (:

OpenStudy (timaashorty):

You're welcome. If someone helped you, don't forget to give them a medal ( which is in the right corner to a comment) to show your appreciation. Also after you got the help you needed close the question. [;

OpenStudy (anonymous):

okayyy